OpenStudy (anonymous):

Calc 3? Describe the geometry of the set points whose coordinates x,y,z satisfy the equation x^2+y^2+z^2+4x+2y-6z-22=0?

OpenStudy (anonymous):

doing quadratic surfaces, are we?

OpenStudy (anonymous):

Try completing the square you will get it in this form (x-a)^2 +(y-b)^2 +(z-c)^2 =k so it's a sphere

OpenStudy (anonymous):

Any ideas?

OpenStudy (anonymous):

Ohh I see I've that before I didn't know if that was what it meant by describing thanks!

OpenStudy (anonymous):

"geometry of the set points" is way of asking what shape

OpenStudy (anonymous):

What happens if it's only like (x-a)^2 +(y-b)^2=K

OpenStudy (anonymous):

then it would be a circle

OpenStudy (anonymous):

what does a and b stand for? the transformation it shifts about? and K is the radius right just making sure I understand it right

OpenStudy (anonymous):

yes , you are right

OpenStudy (anonymous):

okay thanks a lot!

OpenStudy (anonymous):

Calc 3 will get a lot more interesting after this stuff

OpenStudy (anonymous):

Hey sorry I have one more question, it says show that the intersection of that equation with yz is a circle, does that mean that I make x as 0?

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

b/c x=0 on entire y-z plane

OpenStudy (anonymous):

so the equation I got was (x+2)^2 + (y+1)^2 + (z+3)^2 = 34 and I'd make it zero so I'd get just (y+1)^2 + (z+3)^2 = 30

OpenStudy (anonymous):

ohh I see yeah okay thanks

OpenStudy (anonymous):

Okay last thing I swear.... I just want to make sure I'm doing this right b/c I have a test tmr it's just practice

OpenStudy (anonymous):

it says describe the intersection of the obj in part a with object in part c..

OpenStudy (anonymous):

this is part a : (x+2)^2 + (y+1)^2 + (z+3)^2 = 34 and part c (x+2)^2+(y+1)^2=11.. what does that mean ?

OpenStudy (anonymous):

do I combine them? I'm suppose to get 2 circles?

OpenStudy (anonymous):

one is sphere while other is circle

OpenStudy (anonymous):

Yeah that's what I thought but then it says I'm suppose to obtain two circles. So I assumed that intersection meant somehow they were equal

OpenStudy (anonymous):

let me try to plot them

OpenStudy (anonymous):

Can't get it to work if I put them together (x+2)^2 and (y+1)^2 cancels

OpenStudy (anonymous):

you are supposed to obtain two circle?

OpenStudy (anonymous):

It would just be a circle cutting through the sphere

OpenStudy (anonymous):

well you know how if a sphere intersects with a plane you get a circle? I'm assuming if a circle intersects a sphere you get two circles??

OpenStudy (anonymous):

I don't know it says on the sheet, do you have any ideas how to show that?